11. Consider the following set of n-6 observations of x and y given in Table 1 i. Graph the observations in a scatter plot on paper by hand by plotting x on the x-axis and y on the y-axis. Comment on the relationship. Using observations from Table 1, what are the sample means i and-? Using observations from Table 1, calculate the deviations and squared deviations from the sample mean for x: (x 11. iii. -x) and (x )for each i. iv. Using observations from Table 1, compute the following quantities (by hand or using a calculator) and comment on what they represent: a) 721-x)2 (xi-) i-1 (01-9) b)
c) 記に1(xi-x) (yi-y) Write out equation (1), substituting the values you obtained for B1 and Bo in the previous question and interpret your results: v. What does this equation tell us? How could you test whether the population parameter = 0? Draw the above equation (1) as a line into your plot from part (i). Interpret the slope and intercept of the line vi. Table l: Observations on x and y 4 3 10 10
Answer #1

11. (i) The plot is as below.

10 5 10

There seems to be a positive association between x and y, meaning that as x increases, y also tends to increase on average.

(ii) The sample means would be as below.

4+6+ 10+95+3 37 6.1667.

3+8+9+ 10 +6+440 ー 6.666.

(iii) The table is as below. Some values of the mean deviation are adjusted for rounding errors (so that the sum of mean deviation is zero).

i x (x - ar{x}) (x - ar{x})^2
1 4 −2.1667 4.69458889
2 6 −0.1667 0.02778889
3 10 3.8334 14.69495556
4 9 2.8334 8.02815556
5 5 −1.1667 1.36118889
6 3 −3.1667 10.02798889

(iv) (a) Σ(r,-i)2- (4.69458889 +0.02778889+14.69495556+ 8.0281 5556 +1.36118889 + 10.02798889

or (xi -(38.83466668) 6 or Σ(zi-i -6472444447 .

This value represents the variance in x-values, which is a measure of the spread-out of the x-values.

(b) The table is as below. The values of mean deviations of y is adjusted similarly.

i x (x - ar{x}) (y - ar{y}) (x - ar{x})(y - ar{y})
1 4 −2.1667 −3.6666 7.94442222
2 6 −0.1667 1.3333 −0.22226111
3 10 3.8334 2.3333 8.94447222
4 9 2.8334 3.3333 9.44457222
5 5 −1.1667 −0.6667 0.77783889
6 3 −3.1667 −2.6666 8.44432222

i Σ (r,-2)(у,一g) = (7.94442222-0.22226111+8.94447222+9.44457222 6+0.777838898.44432222

or i2 (r,-z)(34-9) -(35.33336666 or (r,-D(y,-9) = 5.888894 44 .

This is the covariance between x and y, which is a measure of the joint variability of x and y. As the covariance is positive, we may confirm our previous comment on the positive association to be correct.

(c) widehat{eta}_1 = rac{rac{1}{n} sum^n_{i=1} (x_i - ar{x})(y_i - ar{y})}{rac{1}{n} sum^n_{i=1} (x_i - ar{x})^2} , and putting the values, we have 5.888894443 6.472444447 or = 0.909840863 .

This is the slope coefficient of the regression, which measures by how much the y-value increase on average for a unit increase in the x value.

(d) widehat{eta}_0 = ar{y} - widehat{eta}_1 ar{x} , and putting the values, we have = 6.6667-0.909840863 * 6. 1667 or = 1.055984 .

(v) The regression equation is as below.

widehat{y}_i = widehat{eta}_0 + widehat{eta}_1 x_i or 1.055984 0.909840863r .

This represents the relation between the average/expected value of y and x values. For x to be zero, the average value of y would be = 1.055984+0.909840863* 0-1.055984 . Also, for a unit change in x, the y would increase by 909840863 units on average.

To test the slope coefficient, we must find the standard error of the slope coefficient. This would be IT มิ , for widehat{sigma}^2 = rac{sum u^2}{n-2} (where u's are residuals, and n-2 is the degree of freedom) or widehat{sigma}^2 = rac{sum (y_i - widehat{y}_i)^2}{n-2} .

Then, we would have ~ ta,n-2 follows the t-distribution with n-2 degree of freedom. To test for eta_1 = 0 , the null hypothesis would be H_0 : eta_1 = 0 and ふチ( , and the t-statistic would be เริ1-0 se(31) se(31) , and if the t > t ,T2-2 , we may reject the null.

(vi) The graph is as below.

10 5 / Regression Line 10.909840863 #beta1 1 beta0 5 10

The intercept value is = 1.055984 , which is where the regression line touches the y-axis, which is average y value when x=0. The slope is = 0.909840863 , which is the increase in average value of y, for a unit increase in x.